OpenStudy (anonymous):

simplify the expression (-6i)(-4i)

zepdrix (zepdrix):

Notice that we simply have 4 things being multiplied together, -6, i, -4, and i. Multiplication is `commutative`, so we can rearrange the order in which we multiply. Let's bring the i's to the back, and combine the numbers first.\[\large\rm (-6)(-4)i\cdot i\]

OpenStudy (anonymous):

24i^2 ???

zepdrix (zepdrix):

So what do you get for the number in front? :)

OpenStudy (anonymous):

is it 24i^2?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we can write this: \[\large \left( { - 6i} \right) \cdot \left( { - 4i} \right) = \left( { - 6} \right) \cdot \left( { - 4} \right) \cdot i \cdot i = 24 \cdot {i^2} = 24 \cdot \left( { - 1} \right) = ...?\]

OpenStudy (anonymous):

is it 24i^2?

OpenStudy (michele_laino):

yes!